An assembly line has 10 stations with times of 1, 2, 3, 4, …, 10, respectively. What is the bottleneck time?
Answer #1

Throughput time= Sum of all the times= 1+2+3+4+5+6+7+8+9+10 = 55

Highest time= 10min

Bottleneck time= Highest time/Throughput time=10/55=0.1818 = 18.18% of the throughput time.